given f(x)=2-2x and g(x)=-2x+2
find f+g

Answer by stanbon(75887)

given f(x)=2-2x and g(x)=-2x+2
find f+g
--------
f+g=2-2x+(-2x+2)
f+g=2-2x-2x+2
f+g=-4x+4
Cheers,
Stan H.